OpenStudy (anonymous):

If 7p + 5q = 2 and 8p - 9q = 17, then -p + 14q =

OpenStudy (anonymous):

I would use substitution to first solve for p and q in the first 2 equations. Then you can plug those in to solve for equation 3

OpenStudy (anonymous):

Using the two equations, find the value of p and q individually and then find the value of -p + 14q. 7p + 5q = 2 7p = 2 - 5q p = (2 - 5q)/7 Substitute this in the second equation to get the value of q, then find p.